OpenStudy (anonymous):

Factor completely: x^2 − 10x + 25 (x − 5)(x − 5) (x + 5)(x + 5) (x + 5)(x − 5) (x − 25)(x − 1)

OpenStudy (anonymous):

umm this answer is first one

OpenStudy (anonymous):

because if u actually solve all four of them the first one only works with the question u gave

OpenStudy (anonymous):

$$x^2-10x+25=x^2-5x-5x+25=x(x-5)-5(x-5)=(x-5)(x-5)$$